OpenStudy (anonymous):

solve the logarithm equation, log[4]x+log4(x+7)=log[4](x+27)

OpenStudy (turingtest):

is that log base 4 ?

OpenStudy (anonymous):

\[\log_{4}x+\log_{4}(x+7)=\log_{4}(x+27) \]

OpenStudy (anonymous):

log[4] x(x+7) = log[4] (x+27) x^2+7x = x + 27 x^2 + 6x - 27 = 0 x^2 + 9x -3x -27 x(x+9)-3(x+9) = 0 (x+9)(x-3) = 0 x = -9, 3 but be ware!! domain : x>0 , x+7>0 so x>-7, x+27>0 so x> -27 "intersection" between them all give you : x>0 so x=3.

OpenStudy (turingtest):

this is the same as \[\log_{4}(x^2+7x)=\log_{4}(x+27)\]because log(a)+log(b)=log(ab)